jagr2713 (jagr2713):

Trick Question #2:If you divide 30 by half and add ten. What do you get?

OpenStudy (anonymous):

25?

jagr2713 (jagr2713):

Nope D:

jagr2713 (jagr2713):

@pooja195 @Kainui

jagr2713 (jagr2713):

30

OpenStudy (anonymous):

why would it be 30?

jagr2713 (jagr2713):

Nope D:

jagr2713 (jagr2713):

no its not 30, the number os 30, read the question

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

70?

jagr2713 (jagr2713):

Correct @Molly_Qian

OpenStudy (anonymous):

"Divide 30 by half" So you have \(\dfrac{30}{\frac{1}{2}} = 2(30)\)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

then you have 60. Add then and there you go.

jagr2713 (jagr2713):

Yep :D

jagr2713 (jagr2713):

Half=0.5

jagr2713 (jagr2713):

divide by 30 and get 60 and add get 70